# Dimitrios M. Thilikos ${}^{,}$

### v3

@@ 4.3 The local structure of a surface
  • For $i\in[c+h],$ $\mathcal{L}_{i}$ is a transaction on $(G_{2},\Omega_{2})$ and $X^{2}_{i}\subseteq X_{2}$ is a segment of $\Omega_{2}.$ The transactions $\mathcal{L}_{1},\dots,\mathcal{L}_{c+h}$ are pairwise disjoint with endpoints in $X_{i}^{2}$ and, moreover, $X_{1}^{2},\dots,X^{2}_{c+h}$ are also pairwise disjoint. For all $i,$ $\mathcal{L}_{i}$ is coterminal with $\mathcal{P}_{2}$ up to level $1$ of $\mathcal{C}_{2}$ in $\rho_{2}$ and $\mathcal{L}_{i}$ is disjoint from $\Sigma_{\delta}(c_{1})$ and disjoint from $\mathcal{P}_{1}.$
  • The transactions $\mathcal{L}_{1},\dots,\mathcal{L}_{h}$ are handle transactions, each of thickness $θ,$ and the transactions $\mathcal{L}_{h+1},\dots,\mathcal{L}_{c+h}$ are crosscap transactions, each of thickness $θ.$
@@
In those circumstances (neglecting the wall $W$) we say that $(W_{1},W_{0},\delta,c_{0},c_{1},\Delta_{0},\mathcal{C}_{1},\mathcal{C}_{2},\mathcal{P}_{1},\mathcal{P}_{2},\gamma)$ is a $\Sigma$-configuration with parameters $(r,|\mathcal{C}_{1}|,|\mathcal{C}_{2}|,|\mathcal{P}_{1}|,θ,c,h).$ The cell $c_{1}$ is called the exceptional cell of $\delta$.
@@
Figure 12: A $\Sigma$-configuration of a graph (a), (b), and (c)) together with a rendition of bounded breadth and depth of its exceptional cell (d)). (a) The vortex-free rendition of the vortex society of $c_{0}$ together with the wall $W_{0}$ which is grounded in this rendition. On the boundary of $c_{0}$ one can find one endpoint of every path in $\mathcal{P}_{1}$. Notice that each such endpoint coincides with a branch vertex of $W_{0}$. (b) A rendition of the complementary society $(G_{2},\Omega_{2})$ to the vortex society of $c_{0}$. This rendition contains the nest $\mathcal{C}_{2}$ of the special cell $c_{2}$, a part of the linkage $\mathcal{P}_{1}$ which connects $W_{0}$ to the boundary of $c_{1}^{\prime}$, the transactions $\mathcal{L}_{i}$ which are either handle or crosscap transactions on $(G_{2},\Omega_{2})$ and reach into the interior of $c_{2}$. The linkage $\mathcal{P}_{2}$ (not highlighted) consists of the subpaths of the members of the $\mathcal{L}_{i}$ connecting the boundary of $c_{1}$ to the boundary of $c_{2}$. Finally, the figure depicts the exceptional cell $c_{1}$. (c) The cylindrical rendition of the vortex society of $c_{1}$, its vortex $c_{1}^{\prime}$, its nest $\mathcal{C}_{1}$, and the remaining part of $\mathcal{P}_{1}$. (d) An alternative rendition of the vortex society of $c_{1}$, this time a rendition of bounded breadth and depth where each vortex is surrounded by a private nest and each such nest together with its vortex lives in a private disk. This rendition is guaranteed to exist due to Subsection 4.3.
@@
Let $G$ be a graph, and let $W$ be an $R$-wall in $G.$ Then either $G$ contains $K_{t}$ as a minor, or there exist integers $c,h\geq 0$ with $c+h\leq{t+1\choose 2},$ a set $A\subseteq V(G)$ of size at most $\mu,$ a surface $\Sigma=\Sigma^{(h,c)}$, and a $\Sigma$-configuration $(W_{1},W_{0},\delta,c_{0},c_{1},\Delta_{0},\mathcal{C}_{1},\mathcal{C}_{2},\mathcal{P}_{1},\mathcal{P}_{2},\gamma)$ of $G-A$ with parameters $(r,M,s_{2},M,M,c,h)$.
Moreover, the vortex society of $c_{1}$ has a rendition $\rho_{1}=(\Gamma_{1},\mathcal{D}_{1})$ of breadth at most $2t^{2}$ and depth at most $\mu$ and such that
  • for each vortex cell $c\in C(\rho_{1})$ there exists a nest $\mathcal{C}_{c}$ of order $M$ in $\rho_{1}$ around the unique disk $\Delta\in\mathcal{D}$ corresponding to $c,$ and
@@
For an illustration of the outcome of Subsection 4.3 see (a), (b), and (d) in Figure 12. Please notice that the structure of the nests for the vortices of $\rho_{1}$ is not made explicit in [32]. However, the existence of these nests follows from the proofs as explained, for example, in [47].
5 The upper bound
We are finally ready to prove the upper bound portion of Theorem 1.2. To do this we will first show how to extract a Dyck-grid of chosen order from a $\Sigma$-configuration. Then, we use this technique to prove a local structure theorem for the case where $\Sigma$ belongs to a fixed closed set of surfaces.
5.1 A local structure theorem for excluding Dyck-grids
@@
Let $c,h,t,r,s_{2}$ be non-negative integers. In case $c=0$ let $c_{0}=h_{0}=0,$ if $c=2p+1$ let $c_{0}=1$ and $h_{0}=p,$ and if $c=2p\neq 0$ let $c_{0}=2$ and $h_{0}=p-1.$ Finally, let $g=2h+c.$ Assume that
@@
If $G$ is a graph with a $\Sigma$-configuration $(W_{1},W_{0},\delta,c_{0},c_{1},\Delta_{0},\mathcal{C}_{1},\mathcal{C}_{2},\mathcal{P}_{1},\mathcal{P}_{2},\gamma)$ with parameters $(r,M,s_{2},M,M,c,h)$ where $\Sigma=\Sigma^{(h,c)}$ then $G$ contains an $\mathscr{D}^{h+h_{0},c_{0}}_{t}$-expansion which is grounded in $\delta$.
@@
In case $c+h=0$ we may conclude immediately as $W_{0}$ is an $r$-wall where $r\geq 16t.$ Hence, $W_{0}$ contains the $(t,4t)$-cylindrical grid, and therefore $\mathscr{D}_{t}^{(0,0)},$ as a minor. Moreover, $W_{0}$ is grounded in $\delta$.
So from now on we may assume that $c+h\geq 1.$ We first collect the cycles necessary for the cylindrical grid which will host the crosscap and handle transactions. Let us number the cycles of $\mathcal{C}_{2}$ as $C_{1},C_{2},\dots,C_{s_{2}}$ as in the definition of a nest. Let $H_{0}$ be the subgraph of $G$ consisting of the union of the cycles $C_{1},\dots,C_{162^{2g}t}.$ For each $i\in[c+h]$ let $Q_{i}$ be the subgraph of $G$ obtained as follows. For each $L\in\mathcal{L}_{i}$ there exist paths $P_{1},P_{2}\in\mathcal{P}_{1}$ such that $L$ is coterminal with $\{P_{1},P_{2}\}$ up to level $162^{2g}t.$ For each $j\in[2],$ let $x_{j}$ be the last vertex of $P_{j}$ on $C_{162^{2g}t}$ when traversing along $P_{j}$ starting on $V(\Omega_{2}).$ Moreover, let $y_{j}$ be the last vertex of $P_{j}$ in $C_{1}\cap L$ when traversing along $P_{j}$ starting on $x_{j}$ and moving towards $\Sigma_{\rho_{1}}(c_{1}^{\prime}).$ Let $L^{\prime}$ be the subpath of $L$ with endpoints $y_{1}$ and $y_{2}.$ Finally, we define $\hat{L}$ to be the path $x_{1}P_{1}y_{1}L^{\prime}y_{2}P_{2}x_{2}.$ Observe, that both $P_{1}$ and $P_{2}$ are orthogonal with $\{C_{1},\dots,C_{162^{2g}t}\},$ and $L^{\prime}$ is internally disjoint from $H_{0}.$ Then $Q_{i}$ is defined as the union of the paths $\hat{L}$ for all $L\in\mathcal{L}_{i}.$ Now, let $H_{1}$ be the union of $H_{0}$ with the graphs $Q_{i}$ for all $i\in[c+h].$
Observe that $H_{1}$ contains a mixed surface grid of order $162^{2g}t$ with $c$ crosscaps and $h$ handles as a minor since each of the linkages $\mathcal{L}_{i}$ has order at least $(162^{2g}+4)t$ which allows forgetting $4t$ of the consecutive $L^{\prime}$-paths.
@@
We now have everything in place to prove the local structure theorem for excluding Dyck-grids. The statement we prove here is more general than the one required Theorem 1.1 as we want to allow for the exclusion of any pair of incomparable surfaces. That is, in the following we fix an Euler-genus $g_{1}=2h_{1}$ and an Euler-genus $g_{2}=c_{2}+2h_{2}$ where $c_{2}\in[2]$ and $g_{1}$ represents a surface of orientable genus $h_{1}$ while $g_{2}$ represents a non-orientable surface. To ensure that the two surfaces are incomparable, we also require $0\leq h_{2}<h_{1}$ since otherwise we could find a Dyck-grid with $h_{1}$ handles as a minor in any Dyck-grid with $h_{2}$ handles and $c_{2}$ crosscaps.
@@
Let $k,h_{1},h_{2},c_{2}$ be non-negative integers where $c_{2}\in[0,2]$ and $0\leq h_{2}<h_{1}.$ Let $g_{1}\coloneqq 2h_{1},$ $g_{2}\coloneqq c_{2}+2h_{2}$ if $c_{2}\neq 0$ and $g_{2}\coloneqq 0$ otherwise, and $g\coloneqq\max\{g_{1},g_{2}\}$. Moreover, fix $t\coloneqq 12(162^{2g}+4)k.$
@@
$\mathbf{g}_{\mathsf{o},g_{1},g_{2}}$ $\coloneqq g_{1}-1$
$\mathbf{g}_{\mathsf{n},g_{1},g_{2}}$ $\coloneqq g_{2}-1\text{ if }c_{0}\neq 0\text{, and }g_{1}\text{ otherwise}$
@@
Moreover, we set
$\mathfrak{D}_{g_{1},g_{2}}\coloneqq\{\mathscr{D}^{(h_{1},0)}=\langle\mathscr{D}^{(h_{1},0)}_{i}\rangle_{i\in\mathbb{N}},\mathscr{D}^{(h_{2},c_{2})}=\langle\mathscr{D}^{(h_{2},c_{2})}_{i}\rangle_{i\in\mathbb{N}}\}$
in case $c_{2}\neq 0$ and
$\mathfrak{D}_{g_{1},g_{2}}\coloneqq\{\mathscr{D}^{(h_{1},0)}=\langle\mathscr{D}^{(h_{1},0)}_{i}\rangle_{i\in\mathbb{N}}\}$
otherwise. Then the local structure theorem $\mathfrak{S}(\mathfrak{D}_{g_{1},g_{2}},\mathbf{a}_{g_{1},g_{2}},\mathbf{g}_{\mathsf{o},g_{1},g_{2}},\mathbf{g}_{\mathsf{n},g_{1},g_{2}},\mathbf{v}_{\mathsf{b},{g_{1},g_{2}}},\mathbf{v}_{\mathsf{d},{g_{1},g_{2}}},\mathbf{R}_{g_{1},g_{2}})$ is valid.
To prove Theorem 5.1 we start by showing that we can always either find a Dyck-grid of order $k$ which corresponds to one of the two fixed Euler-genus $g_{1}$ and $g_{2}$, or we find a $\Sigma$-configuration whose parameters depend only on $k,$ $g_{1},$ and $g_{2}$ where the Euler-genus of $\Sigma$ is strictly less than $g_{1}$ if $\Sigma$ is orientable, and strictly less than $g_{2}$ if $\Sigma$ is non-orientable and $g_{2}\neq 0$ or at most $g_{1}$ if $\Sigma$ is non-orientable but $g_{2}=0,$
@@
Let $k,h_{1},h_{2},c_{2}$ be non-negative integers where $c_{2}\in[0,2]$ and $0\leq h_{2}<h_{1}.$ Let $g_{1}\coloneqq 2h_{1},$ $g_{2}\coloneqq 2h_{2}+c_{2}$ if $c_{2}\neq 0$ and $g_{2}\coloneqq 0$ otherwise, and $g\coloneqq\max\{g_{1},g_{2}\}$. Moreover, let $t\geq 12(162^{2g}+4)k$ be an integer.
Finally, let $r,s_{2},M\geq 1$ be integers with $r\geq 16t,$ $s_{2}\geq 2M\geq 2t,$ and set $\mu\coloneqq\left\lfloor\frac{1}{10^{5}t^{26}}t^{10^{7}t^{26}}\cdot M\right\rfloor.$ Fix
@@
Then either $G$ contains $\mathscr{D}^{(h_{1},0)}_{k}$ or, in case $c_{2}\neq 0,$ $\mathscr{D}^{(h_{2},c_{2})}_{k}$ as a minor, or there exist integers $c,h\geq 0$ with $2h<g_{1},$ $2h+c<g_{2}$ if $c\neq 0$ and $c_{2}\neq 0,$ and $2h+c\leq g_{1}$ if $c\neq 0$ and $c_{2}=0,$ a set $A\subseteq V(G)$ of size at most $\mu,$ a surface $\Sigma=\Sigma^{(h,c)}$, and a $\Sigma$-configuration $(W_{1},W_{0},\delta,o_{0},o_{1},\Delta_{0},\mathcal{C}_{1},\mathcal{C}_{2},\mathcal{P}_{1},\mathcal{P}_{2},\gamma)$ of $G-A$ with parameters $(r,M,s_{2},M,M,c,h)$.
@@
  • for each vortex cell $c\in C(\rho_{1})$ there exists a nest $\mathcal{C}_{c}$ of order $M$ in $\rho_{1}$ around the unique disk $\Delta\in\mathcal{D}$ corresponding to $c,$ and
@@
We use Subsection 4.3 to either obtain a $K_{t}$-minor or a set $A\subseteq V(G)$ of size at most $\mu$ together with integers $c,h\leq{t+1\choose 2},$ a surface $\Sigma$ of Euler-genus $2h+c,$ and a $\Sigma$-configuration $(W_{1},W_{2},\delta,o_{0},o_{1},\Delta_{0},\mathcal{C}_{1},\mathcal{C}_{2},\gamma)$ of $G-A$ with parameters $(r,M,s_{2},M,M,c,h)$ such that the vortex society of $c_{1}$ in $\delta$ has depth at most $\mu$ and breadth at most $2t^{2}$ together with the desired properties for the nests.
In case we find the $K_{t}$-minor, we contain every graph on $12(162^{2g}+4)k$ vertices as a minor and thus, in particular, both $\mathscr{D}^{(h_{1},0)}_{k}$ and $\mathscr{D}^{(h_{2},c_{2})}_{k}.$ So in this case we are done.

### v4

@@ 5 The local structure of a surface
  • For $i\in[\mathsf{c}+\mathsf{h}],$ $\mathcal{L}_{i}$ is a transaction on $(G_{2},\Omega_{2})$ and $X^{2}_{i}\subseteq X_{2}$ is a segment of $\Omega_{2}.$ The transactions $\mathcal{L}_{1},\dots,\mathcal{L}_{\mathsf{c}+\mathsf{h}}$ are pairwise disjoint with endpoints in $X_{i}^{2}$ and, moreover, $X_{1}^{2},\dots,X^{2}_{\mathsf{c}+\mathsf{h}}$ are also pairwise disjoint. For all $i,$ $\mathcal{L}_{i}$ is coterminal with $\mathcal{P}_{2}$ up to level $1$ of $\mathcal{C}_{2}$ in $\rho_{2}$ and $\mathcal{L}_{i}$ is disjoint from $\sigma_{\delta}(c_{1})$ and disjoint from $\mathcal{P}_{1}.$
  • The transactions $\mathcal{L}_{1},\dots,\mathcal{L}_{\mathsf{h}}$ are handle transactions, which are all of thickness $θ,$ and the transactions $\mathcal{L}_{\mathsf{h}+1},\dots,\mathcal{L}_{\mathsf{c}+\mathsf{h}}$ are crosscap transactions, each of thickness $θ.$
@@
In those circumstances (neglecting the wall $W$) we say that $(W_{1},W_{0},\delta,c_{0},c_{1},\Delta_{0},\mathcal{C}_{1},\mathcal{C}_{2},\mathcal{P}_{1},\mathcal{P}_{2},\gamma)$ is a $\Sigma$-configuration with parameters $(r,|\mathcal{C}_{1}|,|\mathcal{C}_{2}|,|\mathcal{P}_{1}|,θ,\mathsf{c},\mathsf{h}).$ The cell $c_{1}$ is called the exceptional cell of $\delta$.
@@
Figure 11: A $\Sigma$-configuration of a graph (a), (b), and (c)) together with a rendition of bounded breadth and depth of its exceptional cell (d)). (a) The vortex-free rendition of the vortex society of $c_{0}$ together with the wall $W_{0}$ which is grounded in this rendition. On the boundary of $c_{0}$ one can find one endpoint of every path in $\mathcal{P}_{1}.$ Notice that each such endpoint coincides with a branch vertex of $W_{0}.$ (b) A rendition of the complementary society $(G_{2},\Omega_{2})$ to the vortex society of $c_{0}.$ This rendition contains the nest $\mathcal{C}_{2}$ of the special cell $c_{2},$ a part of the linkage $\mathcal{P}_{1}$ which connects $W_{0}$ to the boundary of $c_{1}^{\prime},$ the transactions $\mathcal{L}_{i}$ which are either handle or crosscap transactions on $(G_{2},\Omega_{2})$ and reach into the interior of $c_{2}.$ The linkage $\mathcal{P}_{2}$ (not highlighted) consists of the subpaths of the members of the $\mathcal{L}_{i}$ connecting the boundary of $c_{1}$ to the boundary of $c_{2}.$ Finally, the figure depicts the exceptional cell $c_{1}.$ (c) The cylindrical rendition of the vortex society of $c_{1},$ its vortex $c_{1}^{\prime},$ its nest $\mathcal{C}_{1},$ and the remaining part of $\mathcal{P}_{1}.$ (d) An alternative rendition of the vortex society of $c_{1},$ this time a rendition of bounded breadth and depth where each vortex is surrounded by a private nest and each such nest together with its vortex lives in a private disk. This rendition is guaranteed to exist due to Section 5.
Let $G$ be a graph and $W$ be a $n$-wall in $G.$ we say that $G$ has a $K_{t}$-minor grasped by $W$ if there exist connected and pairwise vertex-disjoint subgraphs $X_{1},\dots X_{t}$ such that for all $i\neq j\in[t]$ there exists an edge with one endpoint in $X_{i}$ and the other in $X_{j}$ and for every $i\in[t],$ there exist $t$ distinct rows $R_{x^{i}_{1}},\dots,R_{x^{i}_{t}}$ and columns $C_{y^{i}_{1}},\dots,C_{y^{i}_{t}},$ $x^{i}_{j},y^{i}_{j}\in[n]$ for all $j\in[t]$ such that $R_{x^{i}_{j}}\cap C_{y^{i}_{j}}\subseteq X_{i}$ for all $j\in[t].$
@@
Let $G$ be a graph, and let $W$ be an $R$-wall in $G.$ Then either $G$ contains a $K_{t}$-minor grasped by $W,$ or there exist integers $\mathsf{c},\mathsf{h}\geq 0$ with $\mathsf{c}+\mathsf{h}\leq{t+1\choose 2},$ a set $A\subseteq V(G)$ of size at most $\mu,$ a surface $\Sigma=\Sigma^{(\mathsf{h},\mathsf{c})},$ and a $\Sigma$-configuration $(W_{1},W_{0},\delta,c_{0},c_{1},\Delta_{0},\mathcal{C}_{1},\mathcal{C}_{2},\mathcal{P}_{1},\mathcal{P}_{2},\gamma)$ of $G-A$ with parameters $(r,M,s_{2},M,M,\mathsf{c},\mathsf{h}).$
Moreover, the vortex society of $c_{1}$ has a rendition $\rho_{1}=(\Gamma_{1},\mathscr{D}_{1})$ of breadth at most $2t^{2}$ and depth at most $\mu$ and such that
  • for each vortex cell $c\in C(\rho_{1})$ there exists a nest $\mathcal{C}_{c}$ of order $M$ in $\rho_{1}$ around the unique disk $\Delta\in\mathscr{D}$ corresponding to $c,$ and
@@
Moreover, there exists a function $g\colon\mathbb{N}^{3}\to\mathbb{N}$ and an algorithm that finds one of these outcomes in time $g(t,r,s_{2})|V(G)|^{2},$ where $g(t,r,s_{2})\in 2^{\text{$\mathsf{poly}$}(t+r+s_{2})}.$
For an illustration of the outcome of Section 5 see (a), (b), and (d) in Figure 11. Please notice that the structure of the nests for the vortices of $\rho_{1}$ is not made explicit in [17]. However, the existence of these nests follows from the proofs as explained, for example, in [27].
6 The local structure theorem
We are finally ready to prove the local version, that is a theorem with respect to a large wall, of our structure theorem. To do this, we first show how to extract a Dyck-wall of chosen order from a $\Sigma$-configuration. Then we use this to derive a variant of the local Graph Minors Structure Theorem in the case where we exclude Dyck-grid for some fixed surface as a minor.
6.1 Extracting a Dyck-wall
This section is dedicated to the extraction of a Dyck-wall from a given $\Sigma$-configuration.
@@ 6.1 Extracting a Dyck-wall
Let $\mathsf{c},\mathsf{h},t,r,s_{2}$ be non-negative integers. In case $\mathsf{c}=0$ let $\mathsf{c}_{0}=\mathsf{h}_{0}=0,$ if $\mathsf{c}=2p+1$ let $\mathsf{c}_{0}=1$ and $\mathsf{h}_{0}=p,$ and if $\mathsf{c}=2p\neq 0$ let $\mathsf{c}_{0}=2$ and $\mathsf{h}_{0}=p-1.$ Finally, let $g=2\mathsf{h}+\mathsf{c}.$ Assume that
@@
Let $G$ be a graph that has a $\Sigma$-configuration $(W_{1},W_{0},\delta,c_{0},c_{1},\Delta_{0},\mathcal{C}_{1},\mathcal{C}_{2},\mathcal{P}_{1},\mathcal{P}_{2},\gamma)$ where the parameters are $(r,M,s_{2},M,M,\mathsf{c},\mathsf{h})$ and where $\Sigma=\Sigma^{(\mathsf{h},\mathsf{c})}$ then $G$ contains an $\mathscr{D}^{(\mathsf{h}+\mathsf{h}_{0},\mathsf{c}_{0})}_{t}$-expansion which is grounded in $\delta.$
@@
In case $\mathsf{c}+\mathsf{h}=0$ we may conclude immediately as $W_{0}$ is an $r$-wall where $r\geq 16t.$ Hence, $W_{0}$ contains the $(t,4t)$-cylindrical grid, and therefore $\mathscr{D}_{t}^{(0,0)},$ as a minor. Moreover, $W_{0}$ is grounded in $\delta.$
So from now on we may assume that $\mathsf{c}+\mathsf{h}\geq 1.$ We first collect the cycles necessary for the cylindrical grid which will host the crosscap and handle transactions. Let us number the cycles of $\mathcal{C}_{2}$ as $C_{1},C_{2},\dots,C_{s_{2}}$ as in the definition of a nest. Let $H_{0}$ be the subgraph of $G$ consisting of the union of the cycles $C_{1},\dots,C_{162^{2g}t}.$ For each $i\in[\mathsf{c}+\mathsf{h}]$ let $Q_{i}$ be the subgraph of $G$ obtained as follows. For each $L\in\mathcal{L}_{i}$ there exist paths $P_{1},P_{2}\in\mathcal{P}_{1}$ such that $L$ is coterminal with $\{P_{1},P_{2}\}$ up to level $162^{2g}t.$ For each $j\in[2],$ let $x_{j}$ be the last vertex of $P_{j}$ on $C_{162^{2g}t}$ when traversing along $P_{j}$ starting on $V(\Omega_{2}).$ Moreover, let $y_{j}$ be the last vertex of $P_{j}$ in $C_{1}\cap L$ when traversing along $P_{j}$ starting on $x_{j}$ and moving towards $\sigma_{\rho_{1}}(c_{1}^{\prime}).$ Let $L^{\prime}$ be the subpath of $L$ with endpoints $y_{1}$ and $y_{2}.$ Finally, we define $\hat{L}$ to be the path $x_{1}P_{1}y_{1}L^{\prime}y_{2}P_{2}x_{2}.$ Observe, that both $P_{1}$ and $P_{2}$ are orthogonal with $\{C_{1},\dots,C_{162^{2g}t}\},$ and $L^{\prime}$ is internally disjoint from $H_{0}.$ Then $Q_{i}$ is defined as the union of the paths $\hat{L}$ for all $L\in\mathcal{L}_{i}.$ Now, let $H_{1}$ be the union of $H_{0}$ with the graphs $Q_{i}$ for all $i\in[\mathsf{c}+\mathsf{h}].$
Observe that $H_{1}$ contains a mixed surface grid of order $162^{2g}t$ with $\mathsf{c}$ crosscaps and $\mathsf{h}$ handles as a minor since each of the linkages $\mathcal{L}_{i}$ has order at least $(162^{2g}+4)t$ which allows forgetting $4t$ of the consecutive $L^{\prime}$-paths.
@@
As a direct corollary of Subsection 6.1, if one applies the lemma with parameter $t=2k,$ then inside the $\mathscr{D}_{2k}^{\mathsf{h}+\mathsf{h}_{0},\mathsf{c}_{0}}$-expansion one can find a corresponding Dyck-wall of order $k$ as a subgraph.
We now have everything in place to prove the local structure theorem for excluding Dyck-grids. The statement we prove here is more general than the one required for our main theorem as we want to be able to have more fine-tuned control over the excluded surfaces. That is, in the following we fix an Euler-genus $g_{1}=2\mathsf{h}_{1}$ and an Euler-genus $g_{2}=\mathsf{c}_{2}+2\mathsf{h}_{2}$ where $\mathsf{c}_{2}\in[2]$ and $g_{1}$ represents a surface of orientable genus $\mathsf{h}_{1}$ while $g_{2}$ represents a non-orientable surface. To ensure that the two surfaces are incomparable, we also require $0\leq\mathsf{h}_{2}<\mathsf{h}_{1}$ since otherwise we could find a Dyck-grid with $\mathsf{h}_{1}$ handles as a minor in any Dyck-grid with $\mathsf{h}_{2}$ handles and $\mathsf{c}_{2}$ crosscaps.
@@
Let $k,\mathsf{h}_{1},\mathsf{h}_{2},c_{2}$ be non-negative integers where $\mathsf{c}_{2}\in[0,2]$ and $0\leq\mathsf{h}_{2}<\mathsf{h}_{1}.$ Let $g_{1}\coloneqq 2\mathsf{h}_{1},$ $g_{2}\coloneqq\mathsf{c}_{2}+2\mathsf{h}_{2}$ if $\mathsf{c}_{2}\neq 0$ and $g_{2}\coloneqq 0$ otherwise. Moreover, let $g\coloneqq\max\{g_{1},g_{2}\}$ and fix $t\coloneqq(162^{2g}+4)k.$
@@
Then, given an integer $r\geq t,$ graph $G,$ and an $R$-wall $W$ with $R\geq\mathbf{R}_{g_{1},g_{2}},$ one of the following holds.
  1. 1.
    $G$ contains a (a) $\mathscr{D}^{(\mathsf{h}_{1},0)}_{k}$ or a (b) $\mathscr{D}^{(\mathsf{h}_{2},\mathsf{c}_{2})}_{k}$-expansion $D$ such that $\mathcal{T}_{D}\subseteq\mathcal{T}_{W}$ where case (b) is allowed as an outcome if and only if $g_{2}\neq 0,$ or
  2. 2.
    there exist non-negative integers $\mathsf{c},\mathsf{h}$ where $\mathsf{c}\in[0,2]$ and a surface $\Sigma=\Sigma^{(\mathsf{h},\mathsf{c})}$ where neither $\mathscr{D}^{(\mathsf{h}_{1},0)}_{i}$ nor, in case $\mathsf{c}_{2}\neq 0,$ $\mathscr{D}^{(\mathsf{h}_{2},\mathsf{c}_{2})}_{i}$ embeds for all $i\in\mathbb{N},$ such that there is a set $A\subseteq V(G)$ with $|A|\leq\mathbf{a}_{g_{1},g_{2}},$ a $W$-central $\Sigma$-decomposition $\delta$ of $G-A$ of breadth at most $\mathbf{v}_{\mathsf{b},{g_{1},g_{2}}}$ and depth at most $\mathbf{v}_{\mathsf{d},{g_{1},g_{2}}},$ and an $r$-subwall $W^{\prime}$ of $W-A$ which is flat in $\delta.$
Moreover, there exists a function $g\colon\mathbb{N}^{4}\to\mathbb{N}$ and an algorithm that finds one of these outcomes in time $g(k,\mathsf{h}_{1},r)|V(G)|^{2})$ where $g(k,\mathsf{h}_{1},r)=O(r\cdot 2^{2^{\mathcal{O}(\mathsf{h}_{1})}\text{$\mathsf{poly}$}(k)}).$
To prove Theorem 6.1 we start by showing that we can always either find a Dyck-grid of order $k$ which corresponds to one of the two fixed Euler-genus $g_{1}$ and $g_{2},$ or we find a $\Sigma$-configuration whose parameters depend only on $k,$ $g_{1},$ and $g_{2}$ where the Euler-genus of $\Sigma$ is strictly less than $g_{1}$ if $\Sigma$ is orientable, and strictly less than $g_{2}$ if $\Sigma$ is non-orientable and $g_{2}\neq 0$ or at most $g_{1}$ if $\Sigma$ is non-orientable but $g_{2}=0.$
@@
Let $k,\mathsf{h}_{1},\mathsf{h}_{2},\mathsf{c}_{2}$ be non-negative integers where $\mathsf{c}_{2}\in[0,2]$ and $0\leq\mathsf{h}_{2}<\mathsf{h}_{1}.$ Let $g_{1}\coloneqq 2\mathsf{h}_{1},$ $g_{2}\coloneqq 2\mathsf{h}_{2}+\mathsf{c}_{2}$ if $\mathsf{c}_{2}\neq 0$ and $g_{2}\coloneqq 0$ otherwise, and $g\coloneqq\max\{g_{1},g_{2}\}.$ Moreover, let $t\geq 1$ be an integer.
Finally, let $r,s_{2},M\geq 1$ be integers with $r\geq 16t,$ $s_{2}\geq 2M\geq\max\{2(162^{2g}+4)k,2t\},$ and set $\mu\coloneqq\left\lfloor\frac{1}{10^{5}t^{26}}t^{10^{7}t^{26}}\cdot M\right\rfloor.$ Fix
@@
Then one of the following is true.
  1. 1.
    $G$ contains a $K_{t}$-minor grasped by $W,$
  2. 2.
    $G$ contains an expansion $D$ of (a) $\mathscr{D}^{(\mathsf{h}_{1},0)}_{k}$ or (b) $\mathscr{D}^{(\mathsf{h}_{2},\mathsf{c}_{2})}_{k},$ where outcome (b) is allowed if and only if $\mathsf{c}_{2}\neq 0,$ such that $\mathcal{T}_{D}\subseteq\mathcal{T}_{W},$ or
  3. 3.
    there exist integers $\mathsf{c},\mathsf{h}\geq 0$ and a surface $\Sigma=\Sigma^{(\mathsf{h},\mathsf{c})}$ where neither $\mathscr{D}^{(\mathsf{h}_{1},0)}_{i}$ nor, in case $\mathsf{c}_{2}\neq 0,$ $\mathscr{D}^{(\mathsf{h}_{2},\mathsf{c}_{2})}_{i}$ embeds for all $i\in\mathbb{N},$ such that there is a set $A\subseteq V(G)$ of size at most $\mu$ and a $\Sigma$-configuration $(W_{1},W_{0},\delta,o_{0},o_{1},\Delta_{0},\mathcal{C}_{1},\mathcal{C}_{2},\mathcal{P}_{1},\mathcal{P}_{2},\gamma)$ of $G-A$ with parameters $(r,M,s_{2},M,M,\mathsf{c},\mathsf{h}).$
@@
    • for each vortex cell $c\in C(\rho_{1})$ there exists a nest $\mathcal{C}_{c}$ of order $M$ in $\rho_{1}$ around the unique disk $\Delta\in\mathscr{D}$ corresponding to $c,$ and
@@
Moreover, there exists a function $g\colon\mathbb{N}^{4}\to\mathbb{N}$ and an algorithm that finds one of these outcomes in time $g(k,\mathsf{h}_{1},r)\cdot|V(G)|^{2})$ where $g(k,\mathsf{h}_{1},r)=O(r\cdot 2^{2^{\mathcal{O}(\mathsf{h}_{1})}\text{$\mathsf{poly}$}(k)}).$
@@
We use Section 5 to either obtain a $K_{t}$-minor grasped by $W$ or a set $A\subseteq V(G)$ of size at most $\mu$ together with integers $\mathsf{c},\mathsf{h}\leq{t+1\choose 2},$ a surface $\Sigma$ of Euler-genus $2\mathsf{h}+\mathsf{c},$ and a $\Sigma$-configuration $(W_{1},W_{2},\delta,o_{0},o_{1},\Delta_{0},\mathcal{C}_{1},\mathcal{C}_{2},\gamma)$ of $G-A$ with parameters $(r,M,s_{2},M,M,\mathsf{c},\mathsf{h})$ such that the vortex society of $c_{1}$ in $\delta$ has depth at most $\mu$ and breadth at most $2t^{2}$ together with the desired properties for the nests.
In case we find the $K_{t}$-minor, we are done immediately.